Two building are 300ft apart. from the top of the shorter building the angle of elevation of the top of the taller building is 23 degree, and the angle of depression of the base of the taller building is 36. How tall is each building.
Answer by mananth(16946)

Two building are 300ft apart. from the top of the shorter building the angle of elevation of the top of the taller building is 23 degree, and the angle of depression of the base of the taller building is 36. How tall is each building.
Angle of elevation of taller building from top of small building = 23 deg
.
The extra height of the taller building the distance between the two buildings
form a right triange with the distance of the top of the buildings
tan 23 = height * 300
300 tan 23 = height
300*0.4245 = height
127.35 = height
..
similarly with angle of depression
height = tan 36 * 300
height = 0.7265*300
= 217.95 is the height of the smaller building
height of the taller building = 217.95 + 127.35
=345.30